OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form.

Parth (parthkohli):

Do you know the point-slope form?

OpenStudy (anonymous):

y - y1 = m (x - x1) thats the form

OpenStudy (anonymous):

this is a multi-step problem, first you have to find the slope and then you could use it with whatever set of coordinates out of the two am I right @ParthKohli ?

OpenStudy (anonymous):

can yall help me do it? like i don't understand

OpenStudy (anonymous):

so start out with y2-y1/x2-x1

OpenStudy (anonymous):

after we plug in the points we would get (4-(-4))/(1-(-1)) which would = 4

OpenStudy (anonymous):

4 would equal the slope m so the hard part is done, now we just have to plug it into point slope so the equal would be y - 4 = 4(x-1)

OpenStudy (anonymous):

hope that helps?